a line passing through (-4,4) and (x,-8) is perpendicular to a line with slope 5/12. Find the value of x.
Answer by richwmiller(17219)

new slope is -12/5
y-y/x-x=-12/5
-8-4/x+4=-12/5
-12/x+4=-12/5
x=1